OpenStudy (anonymous):

A projectile is fired vertically upward. Its distance s (in m) above the ground is given by s = 244t - 4.9t2, where t is the time (in s). Find the acceleration of the projectile.

OpenStudy (anonymous):

Is everyone as lost as I was on this one?

OpenStudy (amistre64):

im not

OpenStudy (anonymous):

could you explain please?

OpenStudy (amistre64):

take 2 derivatives

OpenStudy (amistre64):

the derivative of position is velocity, the derivative of velocity is acceleration

OpenStudy (amistre64):

also, assuming this is shot on earth, the only force acting appreciably on a free "falling" object is gravity .... what is the acceleration of gravity?

OpenStudy (anonymous):

Not sure. Doesn't give it in the equation

OpenStudy (amistre64):

oh, it gives it in the equation.

OpenStudy (amistre64):

do you know how to take derivatives?

OpenStudy (anonymous):

Im no expert. We just started yesterday

OpenStudy (amistre64):

well, polynomials are as simple as they get for derivatives; its just an application if what they call a power rule and a constant rule.

OpenStudy (amistre64):

\[D_x[x^n]=x^{n-1}\]\[D_x[k]=0\]

OpenStudy (amistre64):

ugh, forgot to type in the n ... \[D_x[x^n]=n~x^{n-1}\] thats better

OpenStudy (anonymous):

n being the?

OpenStudy (amistre64):

the exponent of course, thats why its tucked up there in the corner of the x .... its the "power" of x .... hence the name, "power rule"

OpenStudy (anonymous):

Got it

OpenStudy (amistre64):

so knowing these properties, how would you attempt a first derivative of the given setup for "s" ?

OpenStudy (anonymous):

f(x) 244t - 4.9t2

OpenStudy (amistre64):

thats the original setup yes, but if you want to refer to it as a function of "x", you might wanna rewrite the ts as xs

OpenStudy (anonymous):

sorry thats not right

OpenStudy (anonymous):

s = 244x - 4.9x2?

OpenStudy (amistre64):

if rewriting the ts as xs makes you more comfortable, thats fine; but we will still need to take a derivative at some point :) 244x will derive to ______ ?

OpenStudy (amistre64):

it might help to see this as:\[244~x^1\]

OpenStudy (anonymous):

I seem to be lost, but would it not derive to 244?

OpenStudy (amistre64):

it would :) and the derivative of a constant is 0, so to recap that term: 244x derives once to 244, and again to 0

OpenStudy (amistre64):

now lets try the second term: -4.9t^2 whats the first derivative?

OpenStudy (anonymous):

Okay I figured I was wrong for getting the same answer :/

OpenStudy (anonymous):

-9.8t?

OpenStudy (amistre64):

correct, and a derivative on that would get us?

OpenStudy (anonymous):

Thats where Im still a little confused. Like prior how you went from 244 to 0

OpenStudy (amistre64):

all we did was take the derivative of a constant: or to put it in power term: 244 x^0 derives to 244(0)x^(-1) = 0

OpenStudy (anonymous):

I got 0 again

OpenStudy (anonymous):

sorry computer froze for a minute

OpenStudy (amistre64):

yeah, the site gets loopy ....

OpenStudy (amistre64):

\[s=244~t^1-4.9~t^2\] \[s'=244~(1)t^0-4.9~(2)t^1\] \[s''=244~(1)(0)t^{-1}-4.9~(2)(1)t^0\]\[s'' = 0-9.8=-9.8\]

OpenStudy (amistre64):

the acceleration on a free falling object is caused by gravity; the acceleration of gravity on the earth is -9.8 \(\frac{meters}{sec^2}\) which means that the force is downward at a magnitude of 9.8

OpenStudy (anonymous):

I see. That helped me see the bigger picture.

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

did it not ask the acceleration upward though?

OpenStudy (anonymous):

thank you for your help!

OpenStudy (amistre64):

no, it asked for the acceleration of the projectile, once the bullet leaves the gun, the only force accelerating it is gravity

OpenStudy (anonymous):

I understand. Thanks again I know it took longer than expected :)

OpenStudy (amistre64):

youre welcome :)